OpenStudy (anonymous):

What polynomial has a graph that passes through the given points? (-2, 2) (-1, -1) (1, 5) (3, 67)

OpenStudy (anonymous):

@DJ♬Øϟϟ¥

OpenStudy (anonymous):

A. y = -x^3 + 4x^2 + 2x - 2 B. y = x^3 + 4x^2 + 2x - 2 C. y = x^4 + 4x^3 + 2x^2 - 2x D. y = x^3 - 4x^2 - 2x - 2

OpenStudy (anonymous):

It's B

OpenStudy (anonymous):

On second thoughts, I'm thinking A. Sorry :(